### v2

# An Empirical Analysis of Tiff’s Impact on American Business Formation

Abstract
This study examines whether the tariff policies, particularly the US-China trade war initiated in 2018, delivered on promises to revitalize American manufacturing and create jobs. Using county-level business application data from 2018-2025, we analyze the relationship between tariff implementation and new business formation through linear regression analysis. Our findings reveal a statistically significant positive association between US tariffs on China and American business applications. However, when Chinese retaliatory tariffs are included in the analysis, their negative coefficient substantially exceeds the positive US tariff effect, suggesting that retaliatory measures largely offset the benefits of protectionist policies. Control variables including inflation rate, federal funds rate, and government spending show significant positive effects on business formation. These results indicate that while protectionist trade policies may stimulate domestic business formation, their effectiveness is significantly diminished by retaliatory responses from trading partners. The study provides evidence that unilateral tariff measures without diplomatic coordination produce limited net benefits, confirming that trade wars create scenarios where potential gains are neutralized by counteractions.
Keywords Business formation, economic revitalization, protectionism, retaliatory tariffs, trade policy impact
1 Introduction
Since 2016, Donald Trump has employed the slogan "Make America Great Again" (MAGA) to rally public support and win his election. One of his most impressive policies promised to his supporters was his tariff policy, particularly his trade war with China since 2018. In his speech, he asserted that his tariffs would lead to a return of American industry and manufacturing jobs. "Jobs and factories will come roaring back into our country," he said while announcing his tariffs (ABC News, 2025).
This raises a fundamental question: Did his policy actually deliver the promised outcomes? In previous research, Fajgelbaum et al. (2020) analyzed the short-term impact of 2018 U.S. tariff policies, finding that tariffs caused significant declines in imports and exports, with complete pass-through of tariffs to U.S. consumers. Similarly, research in the Federal Reserve Board finds that its positive effect from import protection is offset by larger negative effects from rising input costs and retaliatory (Amiti et al., 2019). Moreover, there is plenty of research leading to similar result indicating the effect on various aspects of both the U.S. and China economy, such as Amiti et al. (2020), Amiti et al. (2019), Waugh (2019), Flaaen and Pierce (2019), and Cui and Li (2021).
However, by analyzing the relationship between his tariff policies and business application rates in the United States, we find evidence that contradicts the wholly negative assessments using county-level data from 2018-2025 from U.S. Census Bureau (2025), Federal Reserve Bank of St. Louis (2025), and Peterson Institute for International Economics (2019). Our results suggest that these tariffs may have actually stimulated new business formation, partially fulfilling Trump’s promise of economic revitalization, and provide another view of Trump’s tariffs for further exploration.
2 Data and Methods
In our work, we only consider the data from 2018 to 2025, as the trade war between the U.S. and China began in 2018. This range will contain the full evolution of the trade war, from the initial Section 301 tariffs implemented in July 2018 through subsequent escalations, allowing us to analyze both immediate and medium-term economic impacts.
The most recent data is from April 2025. This 88-month panel provides sufficient variation to examine how tariff shocks propagate through various economic channels over time. The extended timeframe also enables us to control for other macroeconomic factors that might influence the relationship between trade policy and economic outcomes.
We consider the business application for all NAICS from U.S. Census Bureau (2025) as the relevant variable. Meanwhile, the U.S. Tariffs on China from Peterson Institute for International Economics (2019) will be the independent variable we are most interested in. Since the tariff may change multiple times in one month, we will use a weighted sum of the tariff by the number of continuing days every month.
In terms of the control variables we used, we chose the unemployment rate to exclude those necessity entrepreneurs due to the economic shock. We add the inflation rate and the Federal Funds Rate to reflect the effect of the market environment as well. As we also want to isolate the effect of the government force in this case, we use government spending to control it. All of the data above are from Federal Reserve Bank of St. Louis (2025). The data on government spending is only provided in quarterly reports. Thus, we simply divided each by $3$ to fit the monthly data.
Moreover, there is one special control variable we want to mention, which is China’s Tariffs on the U.S. from Peterson Institute for International Economics (2019), as China was the largest trading partner before the trade war. We can also find some fun facts by adding this control variable, and we will discuss it in the following sections.
In order to get the most intuitive result of the tariff, we introduce the simplest linear model to fit the regression analysis through time:
#$\#{Business\_Application}$
$=$ $\beta_{0}+\gamma Z_{US\_Tariffs\_on\_CN}+\beta_{1}X_{unemploy}+\beta_{2}X_{inflation}+\beta_{3}X_{DFF}+\beta_{4}X_{Government}$
$(+\beta_{5}X_{CN\_Tariffs\_on\_US})$
where $\gamma$ refers to the effect of Trump’s tariffs.
One major concern about our data and method may be the range of our data, which includes Biden’s term. However, although Biden didn’t clamor for increasing tariffs or threaten other countries with tariffs, he didn’t decrease Trump’s tariffs either. Thus, we consider including his term to be reasonable and necessary, which is used in other articles like Bown (2022a) and Bown (2022b).
Moreover, since we only want to obtain a rough qualitative analysis, the linear model is enough to provide the relationship between variables. But the value of our model’s coefficient is not very important and can’t be used in the prediction process. The sign and significance level would be the most important term.
3 Result and Discussion
Columns 1-3 and Columns 4-6 in Table 1 report the regression results without and with Chinese tariff variables, respectively. Overall, we find significant relationships between US tariff increases during the trade war and our study variable. Specifically, without considering Chinese tariffs, a one percentage point increase in US tariffs on China is associated with a 1819 unit increase in the dependent variable (Column 1), which is significant at the 1% level. After including Chinese tariffs on the U.S., the coefficient of US tariffs increases to 5142 (Column 4), though its significance level drops to 10%. This decrease in significance level is reasonable, as the correlation between these two variables is quite high.
The positive coefficient of $Z_{US\_Tariffs\_on\_CN}$ provides us with an unexpected result compared with the result by Amiti et al. (2019). Our regression indicates that Trump’s tariffs do have a positive effect on the return of American firms in terms of the number of new firms. However, if we consider the comparison between the two models, the result can be explained as the counterattack by China and other countries offsetting the positive impact of Trump’s tariffs, since the absolute value of $Z_{US\_Tariffs\_on\_CN}$’s coefficient is less than the absolute value of $X_{CN\_Tariffs\_on\_US}$’s coefficient. Then, our result would align with the other article’s result, like Amiti et al. (2019).
Regression w/o Chinese Tariff Regression w Chinese Tariff
Variable Coef. Std. Err. p-value Coef. Std. Err. p-value
Intercept 128400∗∗∗ (40900) 0.002 171200∗∗∗ (55680) 0.003
US Tariffs on CN 1819∗∗∗ (681.7) 0.009 5142 (3019) 0.092
CN Tariffs on US -5296 (4687) 0.262
Unemployment Rate 5563 (5547) 0.319 5958 (5549) 0.286
Inflation 1.226e7∗∗∗ (3.205e6) 0.000 1.266e7∗∗∗ (3.219e6) 0.000
Federal Funds Rate 13610∗∗ (5352) 0.013 13470∗∗ (5344) 0.014
Gov. Spending 64.93∗∗∗ (19.73) 0.001 65.05∗∗∗ (19.70) 0.001
R-squared 0.429 0.438
Adj. R-squared 0.394 0.396
F-statistic 12.33∗∗∗ 10.52∗∗∗
Observations 88 88
Note: p $<$ 0.1, ∗∗ p $<$ 0.05, ∗∗∗ p $<$ 0.01. Standard errors in parentheses.
Table 1: Regression Analysis Results
Among other control variables, inflation rate, federal funds rate, and government spending all show significant positive effects, with inflation rate and government spending being significant at the 1% level and federal funds rate at the 5% level. These results align with macroeconomic theory, which shows that the rise of inflation leads to an increase in investment. Similarly, the higher government spending directly stimulates the government’s power in supporting local firms. While the unemployment rate shows a positive coefficient, it is not statistically significant, which weakly indicates the necessity entrepreneurship. The F-statistics significant at the 1% level, indicating good model fit.
4 Conclusion
This paper examines whether Donald Trump’s tariff policies, particularly the US-China trade war initiated in 2018, delivered on his promise to revitalize American manufacturing and bring back jobs. Through an analysis of US business application data from 2018-2025, the research offers a nuanced perspective on tariff effectiveness. The study reveals that contrary to wholly negative assessments in previous research, Trump’s tariffs demonstrated a statistically significant positive association with new business formation in the United States. However, when Chinese retaliatory tariffs were included in the analysis, the picture became more complex. The absolute value of the Chinese tariff coefficient exceeded that of US tariffs, suggesting that the retaliatory measures may have substantially offset the positive impacts of Trump’s policy.
These findings indicate that protectionist trade policies may stimulate domestic business formation, but their effectiveness is significantly diminished when trading partners implement retaliatory measures. This confirms the economic theory that trade wars create lose-lose scenarios where potential gains from protective tariffs are neutralized by counteractions. For policymakers considering tariffs as a tool for economic revitalization, this research suggests that unilateral protectionist measures without diplomatic coordination are likely to produce limited net benefits due to retaliatory responses.
The study focused primarily on business application rates as a metric for economic revitalization, which captures only one dimension of Trump’s promises. Future research should examine employment figures, wage growth, and manufacturing output to provide a more comprehensive assessment of tariff effectiveness. Additionally, the linear model employed offers a simplified view of complex economic relationships. More sophisticated econometric approaches, including structural models that can better account for industry-specific impacts and regional variations, would enhance our understanding of tariff effects. As the international trade landscape continues to evolve, longitudinal studies tracking the long-term impacts of these tariff policies will be essential for fully evaluating their economic legacy.